### v2

# and symmetric racks

Abstract.
Racks and Legendrian racks are nonassociative algebraic structures based on the framed and Legendrian Reidemeister moves, respectively. Motivated by the classification problem for Legendrian knots, we construct an equivalence of categories between racks and Legendrian racks (and, hence, GL-quandles). We deduce equivalences between kink-involutory racks and Legendrian quandles, involutory racks and Legendrian kei, and the respective pairs of full subcategories whose objects are medial.
As applications, we classify objects in these categories up to order 8 and classify several families of symmetric racks; these results are likely to be of independent interest. In particular, the categories of kei with good involutions, Legendrian kei, and involutory racks are all equivalent.
Key words and phrases:
Classification, equivalence of categories, good involution, involutory rack, kei, Legendrian, medial, quandle, rack, symmetric quandle
2020 Mathematics Subject Classification:
Primary 57K12; Secondary 08A35, 18B99, 20N02
1. Introduction
Legendrian racks are nonassociative algebraic structures used to distinguish Legendrian links in contact three-space. Legendrian racks can be traced back to algebraic structures called kei, which Takasaki [takasaki] introduced in 1942 to study symmetric spaces; quandles, which Joyce [joyce] and Matveev [matveev] independently introduced in 1982 to study links in $\mathbb{R}^{3}$ and $S^{3}$ and conjugation in groups; and racks, which Fenn and Rourke [fenn] introduced in 1992 to study framed links in $3$-manifolds. Kei, quandles, and racks have enjoyed significant study as knot invariants in geometric topology and in their own rights in quantum algebra and group theory.
More recently, various authors have equipped racks with additional structures based on the Legendrian Reidemeister moves. The first work in this direction was by Kulkarni and Prathamesh [original] in 2017. In 2021, Ceniceros et al. [ceniceros] generalized the work of Kulkarni and Prathamesh by introducing Legendrian racks. In turn, Karmakar et al. [karmakar] and Kimura [bi] independently introduced GL-racks, which generalize Legendrian racks, in 2023. In 2025, the author [ta]*Proposition A.2 showed that Legendrian racks can distinguish Legendrian knots not distinguishable by their classical or homological invariants, answering a question of Kimura and reproving a conjecture of Chongchitmate and Ng [atlas].
1.1. Main results
[…]
Theorem 1.1.
There exists an equivalence (actually an isomorphism) of categories $F:\mathsf{LR}\xrightarrow{\sim}\mathsf{Rack}$. Furthermore, $F$ restricts to equivalences (actually isomorphisms) of the following full subcategories:
  1. (1)
    Legendrian quandles $\mathsf{LQ}$ and kink-involutory racks $\mathsf{KIR}$.
  2. (2)
    Trivial Legendrian quandles $\mathsf{LQ}_{\mathrm{triv}}$ and quandles $\mathsf{Qnd}$.
  3. (3)
    Legendrian kei $\mathsf{LK}$ and involutory racks $\mathsf{InvRack}$.
Moreover, if $F$ restricts to an equivalence between two subcategories $\mathcal{C}\subseteq\mathsf{LR}$ and $\mathcal{D}\subseteq\mathsf{Rack}$, then $F$ also restricts to an equivalence of the full subcategories of $\mathcal{C}$ and $\mathcal{D}$ whose objects are medial.
In other words, the novel structures introduced in [ceniceros] are more familiar to the theory than they appear, which is not apparent from the definitions.
Along the way, we obtain several classification results for symmetric racks, motivated by two questions of Taniguchi [symm-class]*Problems 1.1–1.2. These results are likely to be of independent interest. In particular, we show the following.
Theorem 1.2.
The category of symmetric kei is equivalent (actually isomorphic) to the category $\mathsf{LK}$ of Legendrian kei, and similarly for the respective full subcategories whose objects are medial.
1.2. Immediate corollaries
Together, Theorems 1.1 and 1.2 imply the following.
Corollary 1.3.
The category of symmetric kei is equivalent (actually isomorphic) to the category $\mathsf{InvRack}$ of involutory racks, and similarly for the respective full subcategories whose objects are medial.
In tandem with [ta]*Theorem 5.5, Theorem 1.1 also yields the following.
Corollary 1.4.
The categories of Legendrian racks and GL-quandles are equivalent (actually isomorphic), and similarly for the respective full subcategories whose objects are medial.
Figure 1 summarizes the relationships between the various subcategories of $\mathsf{Rack}$ and $\mathsf{LK}$ described in Theorems 1.1 and 1.2 and Corollaries 1.3 and 1.4; cf. [quandlebook]*p. 167.
Two Euler diagrams. The left diagram shows the relationships between various classes of racks. The equivalences between these classes and various classes of Legendrian racks, described in the main theorems, are shown in the right diagram.
Figure 1. Euler diagram showing the relationships between the categories in Theorems 1.1 and 1.2 and Corollaries 1.3 and 1.4. Note that the sizes of the circles are not meant to reflect proportions.
By combining Theorem 1.2 with [ta]*Theorem 4.16, we deduce the categorical center of symmetric kei; cf. Section 2.3.1.
Corollary 1.5.
The center of the category of symmetric kei is $\langle\rho\mid\rho^{2}=1\rangle\cong\mathbb{Z}/2\mathbb{Z}$, the group generated by the collection of all good involutions.
1.2.1. Tabulation
Using GAP [GAP4] and the functors used to prove Theorems 1.1 and 1.2, we classify objects in each of the above categories up to order 8. This uses the classification of GL-racks described in [ta]*Appendix A, which is in turn based on Vojtěchovský and Yang’s [library] library of racks. We provide our code and data in a GitHub repository [my-code]. This work is motivated by the classification problems for GL-racks (see [karmakar]*Section 3) and symmetric racks (see [symm-class]*Problems 1.1–1.2).
Table 1 enumerates our data; cf. [library]*Table 1 and [ta]*Table A.1. By Theorem 1.1, the six rows of Table 1 also count isomorphism classes of racks, medial racks, kink-involutory racks, medial kink-involutory racks, involutory racks, and medial involutory racks up to order 8, respectively. By Theorem 1.2, the last two rows also count isomorphism classes of symmetric kei and medial symmetric kei up to order 8, respectively.
Order 0 1 2 3 4 5 6 7 8
Legendrian racks 1 1 2 6 19 74 353 2080 16023
Medial Legendrian racks 1 1 2 6 18 68 329 1965 15455
Legendrian quandles 1 1 2 5 15 54 240 1306 9477
Medial Legendrian quandles 1 1 2 5 14 48 219 1207 9042
Legendrian kei 1 1 2 5 13 42 180 906 6317
Medial Legendrian kei 1 1 2 5 12 38 168 850 6090
Table 1. Enumeration of various types of Legendrian racks up to order 8, up to isomorphism.
1.3. Structure of the paper
In Section 2, we discuss racks, quandles, medial racks, involutory racks, kei, and a canonical rack automorphism $\theta$ that plays a fundamental role in the theory. We also introduce kink-involutory racks, which are racks for which $\theta$ is an involution.
In Section 3, we discuss several important classes of GL-racks, including Legendrian racks.
In Section 4, we prove Theorem 1.1.
In Section 5, we obtain classification results for symmetric racks. Proposition 5.7 and Corollary 5.8 enhance results of Kamada and Oshiro [symm-quandles-2]*Propositions 3.4 and 3.1. We prove Theorem 1.2 and an extension of the result to involutory racks; see Proposition 5.9. As applications, Corollaries 5.11 and 5.12 strengthen another result of Kamada and Oshiro [symm-quandles-2]*Theorem 3.2 for dihedral quandles.
1.4. Notation
Given a set $X$, we denote the permutation group of $X$ by $S_{X}$, or $S_{n}$ if $|X|=n$. We also denote the composition of functions $\varphi:X\to Y$ and $\psi:Y\to Z$ by $\psi\varphi$.
[…]
We adopt the convention that uses permutations due to its convenience for categorical proofs. One may translate between the two conventions via the formula
$x\triangleright y=s_{y}(x).$
Acknowledgments
I thank Samantha Pezzimenti, Jose Ceniceros, and Peyton Wood for respectively introducing me to Legendrian knot theory, quandles, and GL-racks. I also thank Sam Raskin for advising me during the writing of [ta], the results of which inspired Theorem 1.1.
2. Racks
In this section, we discuss the category $\mathsf{Rack}$ of racks and several important subcategories of $\mathsf{Rack}$. Although we provide all relevant definitions, we also refer the reader to [quandlebook]*Section 5.1 for an accessible introduction to rack theory, [book] for a reference on racks as they concern low-dimensional topology, and [survey] for a survey of modern algebraic literature on racks.
2.1. Racks and quandles
We begin by defining racks and quandles. These algebraic structures are used to construct invariants of framed links and smooth links, respectively.
Definition 2.1.
Let $X$ be a set, let $s:X\to S_{X}$ be a map, and write $s_{x}:=s(x)$ for all elements $x\in X$. We call the pair $(X,s)$ a rack if
$s_{x}s_{y}=s_{s_{x}(y)}s_{x}$
for all $x,y\in X$, in which case we call $s$ a rack structure on $X$. If in addition $s_{x}(x)=x$ for all $x\in X$, then we say that $(X,s)$ is a quandle. We also say that $|X|$ is the order of $(X,s)$.
Example 2.2.
[quandlebook]*Example 99 Let $X$ be a set, and fix $\sigma\in S_{X}$. Define $s:X\rightarrow S_{X}$ by $x\mapsto\sigma$, so that $s_{x}(y)=\sigma(y)$ for all $x,y\in X$. Then $(X,\sigma)_{\mathrm{perm}}:=(X,s)$ is a rack called a permutation rack or constant action rack. (Our notation $(X,\sigma)_{\mathrm{perm}}$ is nonstandard.)
Note that $(X,\sigma)_{\mathrm{perm}}$ is a quandle if and only if $\sigma=\operatorname{id}_{X}$. We call $(X,\operatorname{id}_{X})_{\mathrm{perm}}$ a trivial quandle.
Example 2.3.
[book]*Example 2.13 Let $X$ be a union of conjugacy classes in a group $G$, and define $c^{G}:X\rightarrow S_{X}$ by sending any element $x\in X$ to the conjugation map
$c^{G}_{x}:=[y\mapsto xyx^{-1}].$
Then $\operatorname{Conj}X:=(X,c^{G})$ is a quandle called a conjugation quandle or conjugacy quandle.
Definition 2.4.
Given two racks $R:=(X,s)$ and $(Y,t)$, we say that a map $\varphi:X\to Y$ is a rack homomorphism if
$\varphi s_{x}=t_{\varphi(x)}\varphi$
for all $x\in X$. A rack isomorphism is a bijective rack homomorphism. Endomorphisms and automorphisms of racks are defined in the obvious ways.
We denote the automorphism group of a rack $R$ by $\operatorname{Aut}R$. Finally, the inner automorphism group or right multiplication group of $R$ is the normal subgroup
$\operatorname{Inn}R:=\langle s_{x}\mid x\in X\rangle$
of $\operatorname{Aut}R$.
Example 2.5.
All group homomorphisms $\varphi:G\to H$ are also rack homomorphisms from $\operatorname{Conj}G$ to $\operatorname{Conj}H$. Indeed, for all $x,y\in G$,
$\varphi c^{G}_{x}(y)=\varphi(xyx^{-1})=\varphi(x)\varphi(y)\varphi(x)^{-1}=c^{H}_{\varphi(x)}\varphi(y).$
Example 2.6.
For all racks $(X,s)$, the rack structure $s:X\to S_{X}$ is a rack homomorphism from $(X,s)$ to $\operatorname{Conj}S_{X}$ because, for all $x,y\in X$,
$ss_{x}(y)=s_{s_{x}(y)}=s_{s_{x}(y)}s_{x}s_{x}^{-1}=s_{x}s_{y}s_{x}^{-1}=c^{S_{X}}_{s_{x}}(s_{y})=c^{S_{X}}_{s(x)}s(y).$
2.2. Medial racks
Medial racks are a class of racks notable for their ability to enhance certain invariants of smooth links (see, for example, [Hom]*Example 9 and [enhancements]*Theorems 4.2 and 5.1) and their closed symmetric monoidal structure (see [Hom]*Theorem 12 and [ta]*Theorem 6.7). For references on categorical and algebraic aspects of medial racks, we refer the reader to [rack-roll]*Section 3 and [medial-quandles], respectively.
The following definition analogizes the fact that a group $G$ is abelian if and only if its group operation is a group homomorphism from $G\times G$ to $G$; see [ta]*Section 2.3 for an extended discussion of this analogy.
Definition 2.7.
[rack-roll]*Section 3 Let $R=(X,s)$ be a rack, and consider the product rack $R\times R$. We say that $R$ is medial or abelian if the map $X\times X\to X$ defined by
$(x,y)\mapsto s_{y}(x)$
is a rack homomorphism from $R\times R$ to $R$ [^1] In this paper, we adopt the name “medial” over “abelian.” This is to prevent confusion with commutative racks, which satisfy the much rarer condition that $s_{x}(y)=s_{y}(x)$ for all $x,y\in X$.
Example 2.8.
All permutation racks are medial.
Remark 2.9.
Several equivalent definitions of mediality are well-known, including a pointwise definition and a condition that the rack’s so-called transvection group or displacement group is abelian; see, for example, [stan]*Proposition 2.4. Definition 2.7 will suffice for our purposes.
2.3. The canonical automorphism $\theta_{R}$ of a rack
Every rack $R=(X,s)$ has a canonical automorphism $\theta_{R}$ defined by
$x\mapsto s_{x}(x)$
for all $x\in X$; see [center]*Proposition 2.5. Evidently, $R$ is a quandle if and only if $\theta_{R}=\operatorname{id}_{X}$, so we can loosely think of $\theta_{R}$ as measuring the failure of $R$ to be a quandle. Some authors (e.g., [quandlebook]*p. 149) call $\theta_{R}$ the kink map of $R$ and denote it by $\pi$; this motivates the nomenclature in Section 2.5. When there is no ambiguity, we will suppress the subscript and only write $\theta$ to mean $\theta_{R}$.
Example 2.10.
If $(X,\sigma)_{\mathrm{perm}}$ is a permutation rack, then $\theta=\sigma$.
2.3.1. Centrality of $\theta$
Recall that the center of a category $\mathcal{C}$ is the commutative monoid $Z(\mathcal{C})$ of natural endomorphisms of the identity functor $\mathbf{1}_{\mathcal{C}}$. Concretely, $\eta\in Z(\mathcal{C})$ if and only if, for all objects $R,S$ and morphisms $\varphi:R\to S$ in $\mathcal{C}$, the component $\eta_{R}$ is an endomorphism of $R$, and
$\eta_{S}\varphi=\varphi\eta_{R}.$
For example, if $A\text{-}\mathsf{mod}$ denotes the category of modules over a ring $A$, then the categorical center $Z(A\text{-}\mathsf{mod})$ is isomorphic to the ring-theoretic center $Z(A)$ of $A$.
In 2018, Szymik [center]*Theorem 5.4 showed that $Z(\mathsf{Rack})\cong\mathbb{Z}$ is the free group generated by the collection $\theta$ of canonical automorphisms $\theta_{R}$ for all racks $R$. That is, given any natural endomorphism $\psi:\mathbf{1}_{\mathsf{Rack}}\Rightarrow\mathbf{1}_{\mathsf{Rack}}$ of the identity functor $\mathbf{1}_{\mathsf{Rack}}$, the components $\psi_{R}$ commute with all other rack homomorphisms if and only if there exists an integer $k\in\mathbb{Z}$ such that $\psi_{R}=\theta_{R}^{k}$ for all racks $R$.
2.4. Involutory racks and kei
Important subcategories of $\mathsf{Rack}$ include the category $\mathsf{InvRack}$ of involutory racks and the category of kei or involutory quandles, the latter of which Takasaki [takasaki] introduced in 1943 to study Riemannian symmetric spaces.
More recently, various authors have used involutory racks and kei to construct invariants of unoriented links; see, for example, [quandlebook]*Section 3.1. Another motivation for studying involutory racks and kei comes from the theory of surface-links and the classification problem for symmetric racks; see Section 5. We refer the reader to [quandlebook]*Section 3.1 for a knot-theoretic introduction to kei and [involutory] for a universal-algebraic treatment of involutory racks.
Definition 2.11.
[rack-roll]*Definition 2.3 A rack $(X,s)$ is called involutory if $s_{x}^{2}=\operatorname{id}_{X}$ for all $x\in X$. A kei is an involutory quandle.
Example 2.12.
Given a set $X$ containing more than one element, let $\sigma\in S_{X}$ be any product of disjoint 2-cycles. Then $(X,\sigma)_{\mathrm{perm}}$ is an involutory rack.
Example 2.13.
[quandlebook]*Example 54 Let $A$ be an abelian additive group. Define a rack structure on $A$ by $s_{b}(a):=2b-a$ for all elements $a,b\in A$. Then $T(A):=(A,s)$ is a kei called a Takasaki kei. If $A$ is cyclic, then $T(A)$ is called a dihedral quandle.
2.5. Kink-involutory racks
We introduce kink-involutory racks, a class of racks that includes quandles and involutory racks; cf. Figure 1.
Definition 2.14.
A rack $R=(X,s)$ is called kink-involutory if its canonical automorphism $\theta$ is an involution. Let $\mathsf{KIR}$ be the full subcategory of $\mathsf{Rack}$ whose objects are kink-involutory.
Example 2.15.
Recall that quandles are precisely racks for which $\theta$ is the identity map. It follows that all quandles are kink-involutory. This includes quandles that are not kei, like conjugation quandles of groups $G$ containing an element $g$ such that $g^{2}\notin Z(G)$.
Example 2.16.
For an example of a kink-involutory rack that is neither involutory nor a quandle, consider the rack $R=(X,s)$ with underlying set $X=\{1,2,3,4,5\}$ whose rack structure is given by the permutations
$s_{1}=s_{2}=(12)(345),\quad\quad s_{3}=s_{4}=s_{5}=(12)$
in cycle notation. Then $s_{1}^{2}\neq\operatorname{id}_{X}$, so $R$ is not involutory. Also, $\theta=(12)$ is a nonidentity involution, so $R$ is kink-involutory and not a quandle.
Proposition 2.17.
All involutory racks are kink-involutory.
Proof.
Let $R=(X,s)$ be an involutory rack. For all $x\in X$,
$\theta^{2}(x)=\theta s_{x}(x)=s_{s_{x}(x)}s_{x}(x)=s_{x}s_{x}(x)=x.$
Thus, $\theta^{2}=\operatorname{id}_{X}$. ∎
3. GL-racks and Legendrian racks
In this section, we discuss Legendrian racks and other notable classes of GL-racks. We also apply GL-racks to classification problems for involutory symmetric racks.
3.1. GL-racks
In 2023, Karmakar et al. [karmakar] and Kimura [bi] independently introduced GL-racks to construct invariants of Legendrian links. While the following definition contrasts with the original definitions of Karmakar et al. and Kimura, their equivalence was proven in [ta]*Proposition 3.12.
Definition 3.1.
[ta]*Definition 3.1 Given a rack $R=(X,s)$, a GL-structure on $R$ is a rack automorphism $\mathtt{u}\in\operatorname{Aut}R$ such that $\mathtt{u}s_{x}=s_{x}\mathtt{u}$ for all $x\in X$. We call the pair $(R,\mathtt{u})$ a GL-rack, generalized Legendrian rack, or bi-Legendrian rack.
If in addition $R$ is a quandle or a medial rack, then we also call $(R,\mathtt{u})$ a GL-quandle or medial GL-rack, respectively.
Example 3.2.
[bi]*Example 3.7 Given a permutation rack $P=(X,\sigma)_{\mathrm{perm}}$, a GL-structure on $P$ is precisely a permutation $\mathtt{u}\in S_{X}$ that commutes with $\sigma$.
Definition 3.3.
A GL-rack homomorphism between two GL-racks $(R_{1},\mathtt{u}_{1})$ and $(R_{2},\mathtt{u}_{2})$ is a rack homomorphism $\varphi$ from $R_{1}$ to $R_{2}$ that satisfies
$\varphi\mathtt{u}_{1}=\mathtt{u}_{2}\varphi.$
Let $\mathsf{GLR}$ be the category of GL-racks, and let $\mathsf{GLQ}$ be the full subcategory of $\mathsf{GLR}$ whose objects are GL-quandles.
Remark 3.4.
Virtual racks are algebraic structures used to construct invariants of framed links in certain lens spaces and framed virtual links in thickened surfaces; see, for example, [virtual]*Section 3.2.
By Definition 3.1, GL-racks are precisely virtual racks in which all inner automorphisms $s_{x}$ are endomorphisms of virtual racks. Equivalently, a GL-rack $(X,s,\mathtt{u})$ is a virtual rack in which the operator group of $(X,s)$ identifies $x$ with $\mathtt{u}(x)$ for all $x\in X$; see [fenn]*Section 1.1.
3.1.1. Centrality of GL-structures
By Definition 3.3, all integer powers of GL-structures $\mathtt{u}$ and canonical rack automorphisms $\theta$ lie in the categorical center $Z(\mathsf{GLR})$. In fact, $Z(\mathsf{GLR})\cong\mathbb{Z}^{2}$ is the free abelian group generated by these two collections of automorphisms; see [ta]*Theorem 4.16.
3.2. Legendrian racks
As their name suggests, Legendrian racks are a special class of GL-racks defined below. They were introduced by Ceniceros et al. [ceniceros] in 2021.
Although the following definition differs from the original definition of Ceniceros et al., their equivalence was proven in [ta]*Corollary 3.13.
Definition 3.5.
[ta]*Corollary 3.13 Let $(R,\mathtt{u})$ be a GL-rack. We say that $(R,\mathtt{u})$ is a Legendrian rack if $\theta=\mathtt{u}^{-2}$, in which case we say that $\mathtt{u}$ is a Legendrian structure on $R$.
We say that a Legendrian rack $(R,\mathtt{u})$ is a Legendrian quandle if $R$ is a quandle or, equivalently, if $\mathtt{u}$ is an involution. Similarly, we say that $(R,\mathtt{u})$ is a Legendrian kei if $R$ is a kei.
Let $\mathsf{LR}$, $\mathsf{LQ}$, and $\mathsf{LK}$ be the full subcategories of $\mathsf{GLR}$ whose objects are Legendrian racks, Legendrian quandles, and Legendrian kei, respectively.
Example 3.6.
[bi]*Example 3.6 Let $G$ be a group, and let $z\in Z(G)$ be a central element of $G$. Then multiplication by $z$ defines a GL-structure on the conjugation quandle $\operatorname{Conj}G$. This GL-quandle is a Legendrian quandle if and only if $z^{2}=1$ in $G$.
Example 3.7.
Given any rack $R=(X,s)$, the identity map $\operatorname{id}_{X}$ and the canonical automorphism $\theta_{R}$ are GL-structures on $R$. Also, $R$ is a quandle if and only if $L:=(R,\operatorname{id}_{X})$ is a Legendrian quandle, in which case we say that $L$ is a trivial Legendrian quandle. Let $\mathsf{LQ}_{\mathrm{triv}}$ be the full subcategory of $\mathsf{LQ}$ whose elements are trivial.
4. Proof of Theorem 1.1
4.1. Overview of the proof
We split the proof of Theorem 1.1 into several parts. We work with concrete categories; in this section, all functors will fix morphisms as set maps.
We construct functors $F:\mathsf{LR}\to\mathsf{Rack}$ and $F^{-1}:\mathsf{Rack}\to\mathsf{LR}$ satisfying the following criteria.
Proposition 4.1.
The functors $F$ and $F^{-1}$ are mutually inverse. Moreover, both functors send medial objects to medial objects.
Proposition 4.2.
$F(\mathsf{LQ})=\mathsf{KIR}$, and $F(\mathsf{LQ}_{\mathrm{triv}})=\mathsf{Qnd}$.
Proposition 4.3.
$F(\mathsf{LK})=\mathsf{InvRack}$.
These results will be enough to prove Theorem 1.1.
4.2. Construction of functors
Define $F:\mathsf{LR}\to\mathsf{Rack}$ on objects by
$(X,s,\mathtt{u})\mapsto(X,\mathtt{u}^{3}s),$
where the rack structure $\mathtt{u}^{3}s:X\to S_{X}$ is defined by $x\mapsto\mathtt{u}^{3}s_{x}$.
To construct an inverse functor $F^{-1}:\mathsf{Rack}\to\mathsf{LR}$, define $F^{-1}$ on objects by
$R=(X,s)\mapsto(X,\theta^{-3}_{R}s,\theta_{R}),$
where the rack structure $\theta^{-3}_{R}s:X\to S_{X}$ is defined by $x\mapsto\theta^{-3}_{R}s_{x}$.
Remark 4.4.
By [ta]*Theorem 5.5, there exists an isomorphism of categories ${G:\mathsf{GLQ}\xrightarrow{\sim}\mathsf{Rack}}$ defined by
$(X,s,\mathtt{u})\mapsto(X,\mathtt{u}s),$
and its inverse $G^{-1}:\mathsf{Rack}\xrightarrow{\sim}\mathsf{GLQ}$ is defined by
$R=(X,s)\mapsto(X,\theta^{-1}_{R}s,\theta_{R}).$
Thus, the restriction of $G$ to $\mathsf{LQ}$ is precisely the restriction of $F$ to $\mathsf{LQ}$, and the restriction of $G^{-1}$ to $\mathsf{KIR}$ is precisely the restriction of $F^{-1}$ to $\mathsf{KIR}$.
However, these functors do not agree in general. For example, let $X$ be a set containing at least three elements, let $\sigma\in S_{3}$ be a product of disjoint 3-cycles, and let $R$ be the permutation rack $(X,\sigma)_{\mathrm{perm}}$. Then the underlying racks of $F^{-1}(R)$ and $G^{-1}(R)$ are $R$ and the trivial quandle on $X$, respectively.
4.2.1. Functoriality of $F$ and $F^{-1}$
We briefly verify that $F$ and $F^{-1}$ are functors.
Lemma 4.5.
Let $L=(X,s,\mathtt{u}_{1})$ and $M=(Y,t,\mathtt{u}_{2})$ be Legendrian racks.
  1. (1)
    $F(L)=(X,\mathtt{u}_{1}^{3}s)$ is a rack.
  2. (2)
    If $\varphi:L\to M$ is a GL-rack homomorphism, then $F(\varphi):=\varphi$ is a rack homomorphism from $F(L)=(X,\mathtt{u}^{3}_{1}s)$ to $F(M)=(Y,\mathtt{u}_{2}^{3}t)$.
Hence, $F$ is a covariant functor.
Proof.
The inclusion $\mathtt{u}^{3}\in Z(\mathsf{GLR})$ makes both claims immediate. ∎
Next, we show that $F^{-1}$ is a functor.
Lemma 4.6.
If $R=(X,s)$ is a rack, then $F^{-1}(R)$ is a Legendrian rack.
Proof.
Denote the rack structure of $F^{-1}(R)$ by $t:=\theta^{-3}_{R}s$. Since $\theta^{-3}\in Z(\mathsf{Rack})$, it is straightforward to verify that $R^{\prime}:=(X,t)$ is a rack. Also, the inclusion $\theta\in Z(\mathsf{Rack})$ makes it clear that $\theta_{R}$ is a GL-structure on $R^{\prime}$; that is, $(R^{\prime},\theta_{R})=F^{-1}(R)$ is a GL-rack. In fact, for all $x\in X$,
$\theta_{R^{\prime}}(x)=t_{x}(x)=\theta^{-3}_{R}s_{x}(x)=\theta_{R}^{-2}(x),$
so $F^{-1}(R)$ is a Legendrian rack. ∎
Lemma 4.7.
If $\varphi:R\to S$ is a rack homomorphism, then $F^{-1}(\varphi):=\varphi$ is a GL-rack homomorphism from $F^{-1}(R)$ to $F^{-1}(S)$. Hence, $F^{-1}$ is a covariant functor.
Proof.
This is clear from the inclusions $\theta,\theta^{-3}\in Z(\mathsf{Rack})$. ∎
4.3. Proofs of Propositions 4.14.3
Proof of Proposition 4.1.
First, we show that $FF^{-1}=\mathbf{1}_{\mathsf{Rack}}$ and $F^{-1}F=\mathbf{1}_{\mathsf{LR}}$. Certainly, both compositions fix morphisms. To verify that $FF^{-1}$ fixes objects, let $R=(X,s)$ be a rack. Then
$FF^{-1}(R)=F(X,\theta_{R}^{-3}s,\theta_{R})=(X,s)=R,$
so $FF^{-1}=\mathbf{1}_{\mathsf{Rack}}$. To verify that $F^{-1}F$ fixes objects, let $L=(R,\mathtt{u})$ be a Legendrian rack with $R=(X,s)$, so $\theta_{R}=\mathtt{u}^{-2}$. It follows that
$F^{-1}F(L)=F^{-1}(X,\mathtt{u}^{3}s)=(X,(\mathtt{u}^{3}\theta_{R})^{-3}\mathtt{u}^{3}s,\mathtt{u}^{3}\theta_{R})=(X,s,\mathtt{u})=L,$
so $F^{-1}F=\mathbf{1}_{\mathsf{LR}}$. Hence, $F$ and $F^{-1}$ are mutually inverse.
Finally, the claim that $F$ and $F^{-1}$ send medial objects to medial objects follows straightforwardly from Definition 2.7 and the fact that $\theta^{-3}$ and all GL-structures $\mathtt{u}$ are rack endomorphisms. ∎
Proof of Proposition 4.2.
We verify that $F(\mathsf{LQ})=\mathsf{KIR}$ and $F(\mathsf{LQ}_{\mathrm{triv}})=\mathsf{Qnd}$. Let $L=(X,s,\mathtt{u})$ be a Legendrian quandle. Then $\mathtt{u}^{2}=\operatorname{id}_{X}$, so
$F(L)=(X,\mathtt{u}^{3}s)=(X,\mathtt{u}s).$
For all $x\in X$, the fact that $\mathtt{u}$ is a GL-structure implies that
$\theta^{2}_{F(L)}(x)=\theta_{F(L)}\mathtt{u}s_{x}(x)=\theta_{F(L)}\mathtt{u}(x)=\mathtt{u}^{2}s_{x}(x)=x,$
where we have used the assumption that $(X,s)$ is a quandle. Thus, $F(L)$ is kink-involutory, as desired. Moreover, if $L$ is a trivial Legendrian quandle, then $F(L)=(X,s)$ is the underlying quandle of $L$, as desired.
Conversely, let $R=(X,s)$ be a kink-involutory rack. Then $\theta_{R}^{-2}=\operatorname{id}_{X}$, so
$F^{-1}(R)=(X,\theta_{R}^{-3}s,\theta_{R})=(X,\theta^{-1}_{R}s,\theta_{R}).$
For all $x\in X$,
$\theta^{-1}_{R}s_{x}(x)=\theta^{-1}_{R}\theta_{R}(x)=x,$
so $(X,\theta^{-1}_{R}s)$ is a quandle. Hence, $F^{-1}(R)$ is a Legendrian quandle, as desired. Finally, if $R$ is a quandle, then $\theta_{R}=\operatorname{id}_{X}$, so $F^{-1}(R)=(X,s,\operatorname{id}_{X})$ is a trivial Legendrian quandle. ∎
Proof of Proposition 4.3.
We verify that $F(\mathsf{LK})=\mathsf{InvRack}$. Let $L=(X,s,\mathtt{u})$ be a Legendrian kei, so $\mathtt{u}^{2}=\operatorname{id}_{X}=s_{x}^{2}$ for all $x\in X$. Since $\mathtt{u}$ and $s_{x}$ commute,
$(\mathtt{u}s_{x})^{2}=\mathtt{u}^{2}s_{x}^{2}=\operatorname{id}_{X}.$
Therefore, the rack
$F(L)=(X,\mathtt{u}^{3}s)=(X,\mathtt{u}s)$
is involutory, as desired.
Conversely, if $R=(X,s)$ is an involutory rack, then Proposition 2.17 implies that
$\theta^{-2}_{R}=\operatorname{id}_{X}=s_{x}^{2}$
for all $x\in X$. By Proposition 4.2, $F^{-1}(R)$ is a Legendrian quandle with rack structure $\theta_{R}^{-1}s$. For all $x\in X$,
$(\theta^{-1}_{R}s_{x})^{2}=\theta^{-2}_{R}s_{x}^{2}=\operatorname{id}_{X}$
because $\theta^{-1}\in Z(\mathsf{Rack})$. In other words, the underlying quandle of $F^{-1}(R)$ is involutory, so $F^{-1}(R)$ is a Legendrian kei. This completes the proof of Proposition 4.3 and, hence, the proof of Theorem 1.1. ∎
5. Applications to symmetric rack theory
In this section, we discuss the applications of involutory and Legendrian rack theory to the classification problems for symmetric racks; see, for example, [symm-class]*Problems 1.1–1.2. In particular, we prove Theorem 1.2.
5.1. Symmetric racks
[…]
Definition 5.1.
[symm-quandles]*Definition 2.1 A symmetric rack is a pair $(R,\rho)$ where ${R=(X,s)}$ is a rack and $\rho:X\to X$ is an involution such that
$s_{x}\rho=\rho s_{x}\quad\text{and }s_{\rho(x)}=s^{-1}_{x}$
for all $x\in X$ [^2] Confusingly, some authors (e.g., [rack-roll]) call commutative racks “symmetric.” These two notions are distinct. The map $\rho$ is called a good involution. Note that we do not require $\rho$ to be a rack endomorphism.
Example 5.2.
If $R$ is an involutory rack, then $\theta$ and the identity map are good involutions of $R$; cf. Lemma 5.6. This generalizes the well-known fact that if $R$ is a kei, then the identity map is a good involution of $R$.
Example 5.3.
[symm-quandles]*Example 2.4 Let $G$ be a group. Then the inversion map $g\mapsto g^{-1}$ is a good involution of $\operatorname{Conj}G$.
Definition 5.4.
[symm-quandles]*p. 103 A symmetric rack homomorphism between two symmetric racks $(R_{1},\rho_{1})$ and $(R_{2},\rho_{2})$ is a rack homomorphism $\varphi$ from $R_{1}$ to $R_{2}$ that satisfies
$\varphi\rho_{1}=\rho_{2}\varphi.$
Remark 5.5.
Karmakar et al. [karmakar]*Remark 3.8 observed a striking similarity between the axioms of good involutions and GL-structures. This observation was the inspiration for Theorem 1.2.
5.2. Involutory symmetric racks
Although GL-racks and symmetric racks were introduced independently and for different purposes, we show in this subsection that the former can be used to study the latter. This is motivated by two problems of Taniguchi [symm-class]*Problems 1.1–1.2.
5.2.1. Good involutions are endomorphisms only for involutory racks
First, we state a slight generalization of a result of Kamada and Oshiro [symm-quandles-2]*Proposition 3.4 in the classification of symmetric quandles. Kamada and Oshiro’s original proof applies to our generalization without need for alteration, so we omit it.
Lemma 5.6.
Let $R=(X,s)$ be a rack. Then the following are equivalent:
  1. (1)
    $R$ is involutory.
  2. (2)
    The identity map is a good involution of $R$.
  3. (3)
    $R$ has a good involution $\rho$ that is also a rack automorphism.
  4. (4)
    All good involutions of $R$ are rack automorphisms, and at least one exists.
5.2.2. The group $U_{R}$ of GL-structures on a rack
Motivated by an observation of Karmakar et al. [karmakar]*Remark 3.8, our next goal is to determine a relationship between GL-structures and good involutions of involutory racks.
We begin by recalling the group-theoretic characterization of GL-structures in [ta]*Theorem 4.1. Given a rack $R$, let $U_{R}$ be the set of GL-structures on $R$. It is essentially by definition that $U_{R}$ is the centralizer
$U_{R}=C_{\operatorname{Aut}R}(\operatorname{Inn}R)\trianglelefteq\operatorname{Aut}R.$
Moreover, if $\mathtt{u}_{1},\mathtt{u}_{2}\in U_{R}$, then $(R,\mathtt{u}_{1})\cong(R,\mathtt{u}_{2})$ if and only if $\mathtt{u}_{1}$ and $\mathtt{u}_{2}$ are conjugate in $\operatorname{Aut}R$.
5.2.3. GL-structures and good involutions
The following result is motivated by the problems of classifying symmetric racks both totally and up to isomorphism; see [symm-class]*Problems 1.1–1.2.
Proposition 5.7.
A rack $R$ is involutory if and only if it has a GL-structure that is also a good involution.
In this case, good involutions of $R$ are precisely elements of order 1 or 2 in $U_{R}$, and two good involutions of $R$ yield isomorphic symmetric racks if and only if they are conjugate in $\operatorname{Aut}R$.
Proof.
Let $R=(X,s)$ be a rack, and let $\mathtt{u}\in U_{R}$. If $\mathtt{u}$ is a good involution, then $\mathtt{u}$ satisfies condition (3) of Lemma 5.6, so $R$ is involutory. Conversely, if $R$ is involutory, then by Lemma 5.6, the identity map $\operatorname{id}_{X}$ is a good involution of $R$. Since $\operatorname{id}_{X}\in U_{R}$, the first claim follows.
Now, assume that $R$ is involutory. By part (4) of Lemma 5.6, all good involutions of $R$ are GL-structures. Conversely, if $\mathtt{u}\in U_{R}$, then for all $x\in X$,
$s_{\mathtt{u}(x)}\mathtt{u}=\mathtt{u}s_{x}=s_{x}\mathtt{u}=s_{x}^{-1}\mathtt{u}$
because $\mathtt{u}$ is a rack endomorphism. Since $\mathtt{u}$ is bijective,
$s_{\mathtt{u}(x)}=s^{-1}_{x}.$
Hence, $\mathtt{u}$ is a good involution of $R$ if and only if $\mathtt{u}$ is an involution.
Evidently, an isomorphism of involutory symmetric racks is also an isomorphism of involutory GL-racks; the converse holds if both GL-structures are involutions. This yields the final claim. ∎
5.2.4. Application of Proposition 5.7
Due to Proposition 5.7, combining Example 2.12 with [ta]*Proposition 4.9 yields the following classification of good involutions of involutory permutation racks. This generalizes a result of Kamada and Oshiro [symm-quandles-2]*Proposition 3.1 for trivial quandles.
Corollary 5.8.
Let $X$ be a set, and let $\sigma\in S_{X}$ be a permutation of order 1 or 2 in $S_{X}$. Then the set $\mathcal{I}$ of good involutions of the permutation rack $P=(X,\sigma)_{\mathrm{perm}}$ is precisely the subset of $U_{P}=C_{S_{X}}(\sigma)$ whose elements also have order 1 or 2.
Furthermore, two elements of $\mathcal{I}$ yield isomorphic symmetric racks if and only if they are conjugate in $\operatorname{Aut}P=C_{S_{X}}(\sigma)$.
5.3. Equivalence of symmetric kei and Legendrian kei
In this subsection, we prove Theorem 1.2 and classify good involutions of certain Takasaki kei, including all dihedral quandles.
5.3.1. Discussion of Theorem 1.2
Motivated by a question of Taniguchi [symm-class]*Problem 1.2 about classifying symmetric quandles up to isomorphism, Theorem 1.2 implies that the classification of good involutions of kei is precisely the classification of GL-structures on kei. This result extends the classification of symmetric kei that Kamada and Oshiro [symm-quandles-2]*Proposition 3.4 initiated in 2010.
Proof of Theorem 1.2.
First, we show that the category $\mathsf{LK}$ of Legendrian kei and the category $\mathsf{SK}$ of symmetric kei share exactly the same objects. Let $K=(X,s)$ be a kei, and let $\mathcal{L}\subseteq U_{K}$ be the set of Legendrian structures on $K$. Since $K$ is a quandle, $\mathcal{L}$ is precisely the subset of $U_{K}$ whose elements are involutions. Since $K$ is involutory, it follows from Proposition 5.7 that $\mathcal{L}$ is also the set of good involutions of $K$. This shows that $\mathsf{LK}$ and $\mathsf{SK}$ share exactly the same objects.
By comparing Definitions 3.3 and 5.4, we immediately deduce that the hom-sets in $\mathsf{LK}$ are exactly the same as the corresponding hom-sets in $\mathsf{SK}$. Hence, $\mathsf{LK}$ and $\mathsf{SK}$ are isomorphic. ∎
A similar argument extends Theorem 1.2 from kei to involutory racks.
Proposition 5.9.
The category of involutory symmetric racks is isomorphic to the full subcategory of $\mathsf{GLR}$ whose objects are GL-racks $(R,\mathtt{u})$ where $R$ is involutory and $\mathtt{u}$ is an involution.
Remark 5.10.
If one tries to directly extend the isomorphism in Theorem 1.2 from $\mathsf{LK}$ to the category of involutory Legendrian racks, then the corresponding objects are generally not symmetric racks. For example, let $X=\{1,2,3,4\}$, and let ${\sigma\in S_{4}}$ be the permutation $(13)(24)$ in cycle notation. Then the permutation rack $(X,\sigma)_{\mathrm{perm}}$ is involutory, and the 4-cycle $(1432)\in S_{4}$ is a non-involutory Legendrian structure on $(X,\sigma)_{\mathrm{perm}}$.
Trying to directly extend the isomorphism in Theorem 1.2 from $\mathsf{LK}$ to the category of GL-kei yields a similar outcome. For example, let $n\geq 3$ be an integer, and let $Q$ be a trivial quandle with $n$ elements. Then $Q$ is a kei, but every 3-cycle in $S_{n}$ defines a non-involutory GL-structure on $Q$.
5.3.2. Applications of Theorem 1.2
Using Theorem 1.2, we classify good involutions on infinitely many Takasaki kei. Combining Example 2.13 with [ta]*Example 4.10 and Proposition 4.11 refines the following classification result of Kamada and Oshiro [symm-quandles-2]*Theorem 3.2 for dihedral quandles.
Corollary 5.11.
[symm-quandles-2]*Theorem 3.2 Let $n\geq 2$ be an integer, and define affine transformations $\alpha_{m,b}:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ by $x\mapsto mx+b$. Let $D:=T(\mathbb{Z}/n\mathbb{Z})$ be the dihedral quandle of order $n$.
  1. (1)
    If $n$ is odd, then the only good involution of $D$ is the identity map.
  2. (2)
    If $n=2k$ is even and $k$ is odd, then the only good involutions of $D$ are the identity map and the translation $\alpha_{1,k}$. The corresponding symmetric quandles are nonisomorphic.
  3. (3)
    If $n=2k$ and $k$ is even, then the only good involutions of $D$ are the identity map, the translation $\alpha_{1,k}$, and the affine transformations $\alpha_{k+1,0}$ and $\alpha_{k+1,k}$. Of these good involutions, only the last two yield isomorphic symmetric quandles.
Similarly, [ta]*Proposition 4.9 allows us to generalize the first part of [symm-quandles-2]*Theorem 3.2.
Corollary 5.12.
If $A$ is an abelian group without 2-torsion, then the only good involution of the Takasaki kei $T(A)$ is the identity map.